OpenStudy (anonymous):

Will FAN! What is the value of a? 6(A+2) +3 = 3(10-3a) a =_____

OpenStudy (anonymous):

A=1 would you like me to explain?

OpenStudy (anonymous):

Yea sure!

OpenStudy (anonymous):

first you need to distribute. 6A+12+3=30-9A

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then you need to add like terms. 6A+15=30-9A next you get A alone on one side by subtracting the -9A and 15 from each side. 15A=15

OpenStudy (anonymous):

then you divide the 15 out of the A A=1

OpenStudy (anonymous):

Oh ok. I understand thank you! :)

OpenStudy (anonymous):

yw.